OpenStudy (anonymous):

integral 1/2 to sqrt 3/2 6/sqrt of 1-t^2

OpenStudy (anonymous):

is this \[6\int_{\frac{1}{2}}^{\frac{\sqrt{3}}{2}}\frac{1}{\sqrt{1-t^2}}dt\]?

OpenStudy (anonymous):

anti derivative of \[\frac{1}{\sqrt{1-t^2}}\] is \[\sin^{-1}(x)\] so you can evaluate at upper limit, lower limit, subtract, and then multiply the result by 6